MOA-II Obsevation Log Monthly Summary 2007/5

Return to index

Previous month Next month

TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB6230-70828524866289180
LMC1492-155563351101500
SMC365-384203910900
ALERT-000
GRB-000
TOTAL---9355256401580 (111h33m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
199188195238794158165161770157154149
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
2201441871611701401341291271260


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
161615151419545351211918
Field13Field14Field15Field16
202000


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
20190000000000


SPECIAL TARGET NAME

CategoryTargetNrun


CCD TEMPERATURE Previous month Next month




VACUUM Previous month Next month




TELESCOPE TEMPERATURE Previous month Next month




COMMENTS
2007/5/1: no observation / night time = 42900(sec)

2007/5/2: No observation / night time = 43020(sec)

2007/5/3: D2236,D2237(60sec) and D2238(30sec) has some noise in chip1,2 D2239(300sec) have some noise in chip1 / night time = 43140(sec)

2007/5/4: full moon is close bulge , D2240(60sec) is OK / night time = 43320(sec)

2007/5/5: moon is close bulge , D2241(20sec) is OK / night time = 43440(sec)

2007/5/6: moon is colse to bulge / night time = 43560(sec)

2007/5/7: D2245(60sec) chip1 have noise / night time = 43620(sec)

2007/5/8: D2246(60sec) , D2247(300sec) / night time = 43740(sec)

2007/5/9: D2248(300sec) / night time = 43860(sec)

2007/5/10: D2249-60: OK. D2250-300: OK. Dark noise for a vacume sensor (hot cathod + pyrani) was tested at D2251-D2253. The current broken vacuume sensor was replaced by a repaired one. / night time = 43980(sec)

2007/5/11: Dark: D2254 (300sec) OK. , D2255 (60sec) OK. / night time = 44100(sec)

2007/5/12: Dark: D2256 (60sec) OK. / night time = 44220(sec)

2007/5/13: Dark: D2258 (60sec) and D2259 (300sec) OK. / night time = 44340(sec)

2007/5/14: B6559 was taken by V-filter (240sec). Dark: D2260 (60sec) OK, D2262 (240sec) OK. / night time = 44460(sec)

2007/5/15: Today, winds come suddenly. Winds >10m/s sometimes occur suddenly after 1 hour of no wind. Dark: D2263 (300sec) and D2266 (60sec). / night time = 44580(sec)

2007/5/16: Pumping started from 4:00am / night time = 44700(sec)

2007/5/17: Dark: D2267-60-a,b,c, D2268-300-a.b, D2269-300-a OK. (D2268-300-c: only chip 7 has a problem). GB16 was exposed for MOA-gb174. / night time = 44820(sec)

2007/5/18: B6647: I band, 120 sec. Dark: D2270 (60sec), D2271 (120sec), D2272 (300sec) / night time = 44880(sec)

2007/5/19: Drk: D2273 (60sec), D2274 (300sec) OK. / night time = 44940(sec)

2007/5/20: Dark: D2275 OK. / night time = 45060(sec)

2007/5/21: Dark: D2276 (60sec), D2277 (300sec): OK / night time = 45180(sec)

2007/5/22: Dark: D2280 / night time = 45240(sec)

2007/5/23: Dark: D2285 OK. / night time = 45360(sec)

2007/5/24: Dark: D2286 (60sec), D2287 (300sec) OK. / night time = 45420(sec)

2007/5/25: D2296(60sec) is OK / night time = 45480(sec)

2007/5/26: No observation. D2301(60sec) is OK. / night time = 45600(sec)

2007/5/27: D2302(60sec), D2303(10sec), D2304(20sec) and D2305(300sec) are OK. / night time = 45720(sec)

2007/5/28: D2306(60sec), D2307(100sec) and D2308(300sec) are OK. / night time = 45720(sec)

2007/5/29: D2309(60sec) is OK. / night time = 45780(sec)

2007/5/30: D2310(10sec), D2311(20sec), D2312(40sec), D2313(60sec) and D2314(300sec) are OK. Almost all GB images are cloudy. Be careful when you analyse. / night time = 45900(sec)

2007/5/31: D2317(20sec) and D2321(60sec) are OK. / night time = 45900(sec)

TOTAL NIGHT TIME: 1381980(sec)

OPERATING RATE: 44.3 %
GB: 35.0 %
LMC: 8.4 %
SMC: 0.9 %
ALERT: 0.0 %
GRB: 0.0 %

Return to index

Previous month Next month